OpenStudy (anonymous):

The equation x2 + y2 − 2x + 2y − 1 = 0 is the general form of the equation of a circle. What is the standard form of the equation? (x − 1)2 − (y − 1)2 = 1 (x − 1)2 + (y + 1)2 = 3 (x + 1)2 + (y − 1)2 = 1 (x − 1)2 + (y + 1)2 = 1

OpenStudy (foreverandalways13):

@Luigi0210

OpenStudy (anonymous):

@geerky42

geerky42 (geerky42):

Arrange equation: \[x^2 - 2x + y^2 + 2y = 1\] Now we need to insert something here: \[(x^2 - 2x + \underline{\text{ ? }} )+ (y^2 + 2y + \underline{\text{ ? }}) = 1\] So we can complete the square it. Any idea?

geerky42 (geerky42):

Saying we have something like \(a^2 + b~a\). To complete the square, we should add \(\left(\frac{b}{2}\right)^2\) to both sides, so we have \(a^2 + b~a+\left(\frac{b}{2}\right)^2 \Rightarrow\left(a+\frac{b}{2} \right)^2\)

OpenStudy (anonymous):

@mathmath333

OpenStudy (anonymous):

@peachpi